### v2

# Odd-parity Magnetism Driven by Antiferromagnetic Exchange

Abstract
Realizing odd-parity, time-reversal-preserving, non-relativistic spin splitting is a central goal for spintronics applications. We propose a group-theory-based microscopic framework to induce odd-parity spin splitting from coplanar antiferromagnetic (AFM) states without spin-orbit coupling (SOC). We develop phenomenological models for 421 conventional period-doubling AFM systems in non-symmorphic space groups and construct minimal microscopic models for 119 of these. We find that these AFM states can attain three possible competing ground states. These ground states all break symmetries in addition to those broken by the usual AFM order. Specifically, they give rise to either odd-parity spin-splitting, nematic order, or scalar odd-parity order related to multiferroicity. Our microscopic theories reveal that the odd-parity spin-splitting energy scale is generically large and further reveal that the scalar odd-parity order gives a non-zero Berry curvature dipole without SOC. We identify 67 materials in the Magndata database for which our theory applies. We provide DFT calculations on Fe-based materials that reveal an $h$-wave spin splitting consistent with our symmetry arguments and apply our microscopic model to determine the non-relativistic Edelstein response for CeNiAsO.
Introduction: Generating spin-split band structures without spin-orbit coupling has emerged as a central goal in quantum magnetism. This interest is driven by the discovery of altermagnets, a newly classified collinear compensated magnetic state that breaks time-reversal symmetry Hayami et al. (2019); Yuan et al. (2020); Mazin et al. (2021); Šmejkal et al. (2022, 2022); Bhowal and Spaldin (2024). Altermagnets feature a large non-relativistic even-parity spin-splitting Krempaský et al. (2024) that enable spin-transport González-Hernández et al. (2021); Han et al. (2024) and spin calitronics Cui et al. (2023); Liu et al. (2024). Recently, altermagnetism has inspired generalizations to coplanar magnetic states which allow large odd-parity time-reversal preserving non-relativistic spin-splittings Birk Hellenes et al. (2023); Brekke et al. (2024); Chen et al. (2024); Xiao et al. (2024); Jiang et al. (2024). These $p$-wave, or more generally, odd-parity magnets exhibit spin-splittings that are odd under a sign change of the momentum. They provide nonrelativistic variants of the familiar Rashba and Dresselhaus SOC and hence are likely to play an important role in spintronics Manchon et al. (2015).
Spontaneous odd-parity time-reversal preserving transitions in the spin sector have previously been considered as Pomeranchuk instabilities of a spherical Fermi surface Wu et al. (2007); Kiselev et al. (2017); Wu et al. (2018). Specifically, it was suggested that a $p$-wave spin splitting (or $p$-wave magnetism) spontaneously emerges from an $l=1$ orbital angular momentum Pomeranchuk instability in the spin channel Wu et al. (2007). However, it was later shown that this instability is forbidden due to local spin conservation Kiselev et al. (2017), rendering this mechanism for a non-relativistic $p$-wave spin-splitting ineffective. In the more recent analysis of $p$-wave magnetism Birk Hellenes et al. (2023); Brekke et al. (2024); Chen et al. (2024); Xiao et al. (2024); Jiang et al. (2024), this no-go theorem is implicitly circumvented by inducing $p$-wave magnetism from coplanar antiferromagnets (AFMs). Since such $p$-wave magnetic order is a secondary order parameter, the energetic arguments of Ref. Kiselev et al. (2017), which rely on $p$-magnetism being a primary order parameter, no longer apply. This opens a viable route to generating non-relativistic odd-parity magnetism.
Although significant progress has been made in applying spin-space group symmetry arguments to induce odd-parity time-reversal preserving magnetism Birk Hellenes et al. (2023); Chen et al. (2024); Xiao et al. (2024); Jiang et al. (2024), realistic microscopic models that allow for deeper insight into the properties and origins of this state are rare Brekke et al. (2024). Here we present a group-theory-based microscopic template for inducing odd-parity spin splittings from 421 AFMs and provide explicit microscopic theories for over 100 of these. This provides realistic microscopic models for $p$-wave, $f$-wave, and $h$-wave spin splittings. We show that these spin splittings are naturally induced by coplanar AFM states when no SOC is present. Furthermore, we find that the AFM states we consider here generically exhibit some sort of translation invariant induced order. Specifically, we find that either non-relativistic odd-parity spin-splitting, nematicity, or scalar odd-parity order must emerge in the AFM state. The nematicity that emerges generalizes the AFM-driven nematic order discussed in Fe-based AFMs Kreisel et al. (2020); Christensen et al. (2019). The scalar odd-parity order that emerges appears in the commensurate AFM state of orthorhombic RMnO3 rare earth manganites that exhibit improper ferroelectricity Kimura et al. (2003); Hur et al. (2004); Sergienko et al. (2006); Cheong and Mostovoy (2007).
In the following, we first develop a phenomenological theory for this class of AFM states, identify relevant material candidates in the Magndata database, and then turn to general symmetry-based microscopic theories.
Induced odd-parity magnetism: Here, two key ingredients are required to induce odd-parity spin splittings. First, the AFM state must belong to an irreducible representation (IR) that is at least two-dimensional (2D) so that there are two magnetic degrees of freedom. Second, the inversion operator for this IR must be non-trivial. Both ingredients are natural consequences of non-symmorphic space groups.
These two symmetry ingredients are illustrated in the simplest unit cell doubling AFM in Fig. 1. The nonmagnetic unit cell consists of two sublattice atoms. The two sublattice magnetic moments, here labeled ${\bf S_{1}}$ and ${\bf S_{2}}$, typically form a 2D IR for AFM in non-symmorphic space groups. As described below, consideration of the inversion operator naturally leads to two classes. In the class 1 (top panel), the two atoms are related by inversion, whereas in the class 2 (bottom panel), each atom sits at an inversion center.
Here ${\bf S_{1}}\times{\bf S_{2}}$ is an important quantity, as we discuss later, it generates the odd-parity spin-splitting. This induced order is time-reversal invariant, as time-reversal symmetry flips both spins, and is translationally invariant. The inversion symmetry acts non-trivially on ${\bf S_{1}}$ and ${\bf S_{2}}$ (into blue dashed arrows). In the class 1, inversion interchanges ${\bf S_{1}}$ and ${\bf S_{2}}$. In the class 2, inversion leaves ${\bf S_{1}}$ unchanged but flips ${\bf S_{2}}$. In both cases, the resulting spin splitting ${\bf S_{1}}\times{\bf S_{2}}$ is odd under inversion. Here ${\bf S_{1}}\times{\bf S_{2}}$ can also be interpreted as a non-relativistic version of the Dzyaloshinsky-Moriya vector Dzyaloshinsky (1958); Moriya (1960).
Figure 1: Spin configuration of the coplanar AFM state, denoted by red solid arrows. Its inversion pair is illustrated by blue dashed arrows. The inversion center is at the X mark. Black solid lines denote the nonmagnetic unit cell. (Top) For sites related by inversion. (Bottom) For sites at the inversion center.
It is possible to identify all space groups and AFM ordering wavevectors that give rise to the two required symmetry ingredients. The nontrivial inversion operator requires non-symmorphic space groups, where inversion does not commute with all symmetries at certain time-reversal and inversion invariant momenta (TRIM). This includes 325 IRs for $p$-wave, 84 IRs for $f$-wave, and 12 IRs for $h$-wave spin splittings. The list of space groups and wavevectors is in SM. In these 421 distinct scenarios, the AFM state with odd-parity magnetism can be stabilized through a single continuous phase transition. These 2D IRs are not restricted to cases with only two atoms per nonmagnetic unit cell. For systems with more atoms, the theory remains applicable. However, $\bf S_{1}$ (and $\bf S_{2}$) now describes spin order across multiple atoms in the non-magnetic unit cell. The $p$-, $f$-, and $h$-wave spin splittings are identified by analyzing other symmetry operators at these TRIMs.
Figure 2: (Left) FeSe coplanar magnetic state (Center) DFT band dispersion along the momentum cut ±$\pm(\pi,0.4\pi,0.5\pi)$. (Right) Fermi surface ($k_{z}=0.4\pi$ section) exhibiting $h$-wave spin splitting. The band spin induced is represented by a red-white-blue color code.
To test our symmetry-based arguments, we have examined if the predicted $h$-wave $k_{x}k_{y}k_{z}(k_{x}^{2}-k_{y}^{2})\sigma_{z}$ spin-splitting appears within a DFT calculation for FeSe with coplanar order at the M-point $(\pi,\pi,0)$. Collinear AFM ordering at the M point is known to generate the nematic order observed in many Fe-based materials Kreisel et al. (2020). A coplanar magnetic order is sometimes observed in this class of materials Ding et al. (2017); Stadel et al. (2022). In particular, this coplanar state appears in LaFeAs1-xPx Stadel et al. (2022) with space group 129, providing a materials realization of this state. As shown in Fig. 2, the DFT results for this coplanar state agree with our symmetry prediction. Furthermore, this calculation reveals that the spin-splitting, even for a highly nodal $h$-wave state, can be as large as 0.1 eV.
Phenomenological model: We apply Landau theory to describe the 421 different AFM phase transitions and analyze all the possible induced orders. In the absence of SOC, the free energy density of the 2D IR is Fernandes et al. (2016)
$\begin{split}f&=\alpha(T)\sum_{i}{\bf S}_{i}\cdot{\bf S}_{i}+\beta_{1}(\sum_{i}{\bf S}_{i}\cdot{\bf S}_{i})^{2}\\ &+\beta_{2}({\bf S}_{1}\cdot{\bf S}_{2})^{2}+\beta_{3}({\bf S}_{1}\cdot{\bf S}_{1})({\bf S}_{2}\cdot{\bf S}_{2}).\end{split}$ (1)
Here, we have included all the symmetry-allowed terms to quartic order in $\bf S_{1,2}$. The continuous phase transition occurs as $\alpha(T)\rightarrow 0$. Depending on $\beta_{i}$, there are three competing AFM states with different uniform induced orders (Fig. 3).
Figure 3: Three competing phases and their induced orders. Phase boundaries are at $\beta_{3}=0$, $\beta_{2}=0$, and $\beta_{2}+\beta_{3}=0$.
Among these induced orders, the spin-vector ${\bf S}_{1}\times{\bf S}_{2}$ exhibits odd-parity spin splitting, while the other two are spin-scalars, whose characteristics differ depending on the site symmetry. In the class 1, $(|{\bf S}_{1}|^{2}-|{\bf S}_{2}|^{2})$ is odd-parity, while $({\bf S}_{1}\cdot{\bf S}_{2})$ is even-parity (nematicity). In the class 2, their physical meanings are opposite. This phenomenological model describes 45 odd-parity states (14 spin-vector, 24 spin-scalar, and 7 mixed by SOC) and 22 nematic materials in the Magndata database Gallego et al. (2016) (list included in SM).
The odd-parity spin order ${\bf S_{1}}\times{\bf S_{2}}$, as a spin vector, can generate non-relativistic spin-related odd-parity responses, including odd-parity spin-splitting and the non-relativistic Edelstein effect. On the other hand, the scalar odd-parity orders, which are invariant under spin rotation, produce non-relativistic spin-independent odd-parity responses, including nonlinear I-V transport and ferroelectric-like order. To estimate the magnitudes of the induced orders and to evaluate the response properties of the magnetic states, we develop symmetry-based minimal microscopic models.
Microscopic models: For pedagogical reasons, we first introduce a simplified model. Specifically, we consider the class 1 shown in Fig. 1 and keep only NN hopping $t$ and the magnetic order $\vec{J}_{i}$
$H=\begin{pmatrix}\epsilon_{k}+t\cos\frac{k_{x}}{2}\tau_{x}&O_{M}\\ O_{M}^{\dagger}&\epsilon_{k+Q}-t\sin\frac{k_{x}}{2}\tau_{y}\end{pmatrix}.$ (2)
Here $\tau_{i}$ are Pauli matrices for nonmagnetic sublattices and the on-site exchange driven spin ordering is described by $O_{M}=\vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma}$. As in the usual SDW theoryAuerbach (2012), $H$ features sublattice-independent hopping $\epsilon_{k}$ and $\epsilon_{k+Q}$ on the diagonal blocks. The model also exhibits inter-sublattice hopping $t_{k}\tau_{x}$ and $t_{k+q}\tau_{y}$ on the diagonal ($t_{k}=t\cos\frac{k_{x}}{2}$). The transformation from $\tau_{x}$ to $\tau_{y}$ is a consequence of our choice of Fourier basis. In the nonmagnetic unit cells, sites $a_{2}$ and $b_{2}$ are equivalent, and their hopping to site $a_{1}$ can give the real term $t(e^{ik_{x}/2}+e^{-ik_{x}/2})\tau_{x}=t\cos\frac{k_{x}}{2}\tau_{x}$. In the doubled unit cell, since $a_{2}$ and $b_{2}$ are inequivalent, these two hoppings are described by different operators, resulting in both real ($t\cos\frac{k_{x}}{2}\tau_{x}$) and imaginary ($t\sin\frac{k_{x}}{2}\tau_{y}$) terms. The transformation from $\tau_{x}$ to $\tau_{y}$ can also be understood as a consequence of time-reversal symmetry that exists in the nonmagnetic state: the odd time-reversal symmetry of $\sin\frac{k_{x}}{2}$ can only be compensated by transforming $\tau_{x}$ to $\tau_{y}$.
The $8\times 8$ pedagogical Hamiltonian Eq. (2) captures the essence of nonsymmorphic crystal symmetries with the two sublattices, translational symmetry breaking with unit cell doubling, and spin orderings. We now extend the model to general tight-binding models with Wyckoff position of multiplicity two, including all symmetry-allowed hopping for different nonsymmorphic space groups. This minimal model captures 13 nonsymmorphic materials (listed in SM) with odd-parity AFM. The models for the two inversion classes differ slightly. We explicitly consider the class 1 below and include the models for the class 2 in the SM. The Hamiltonian is
$\begin{split}&H({\bf k})=\begin{pmatrix}h_{0}({\bf k})&O_{M}\\ O_{M}^{\dagger}&\widetilde{h_{0}}({\bf k+Q})\end{pmatrix}\\ &=\begin{pmatrix}\epsilon_{0}+t_{0}+t_{x}\tau_{x}+t_{y}\tau_{y}&\vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma}\\ \vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma}&\epsilon_{0}-t_{0}+\widetilde{t_{x}}\tau_{x}+\widetilde{t_{y}}\tau_{y}\end{pmatrix}\\ &\equiv\epsilon_{0}+t_{1}\rho_{0}\tau_{x}+t_{2}\rho_{0}\tau_{y}+t_{3}\rho_{z}\tau_{x}+t_{4}\rho_{z}\tau_{y}\\ &+t_{0}\rho_{z}+\rho_{x}(\vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma}),\end{split}$ (3)
where $t_{1}\equiv(t_{x}+\widetilde{t_{x}})/2$, $t_{2}\equiv(t_{y}+\widetilde{t_{y}})/2$, $t_{3}\equiv(t_{x}-\widetilde{t_{x}})/2$, and $t_{4}\equiv(t_{y}-\widetilde{t_{y}})/2$. The Pauli matrices $\tau_{i}$ and $\rho_{i}$ act in sublattices $(1,2)$ and $(a,b)$, respectively. The time-reversal operator and inversion operator are $T=i\sigma_{y}K$ and $I=\tau_{x}$, with momentum flip. The first block, $h_{0}({\bf k})$ is the same as the nonmagnetic Hamiltonian, containing k-even $\tau_{x}$ and k-odd $\tau_{y}$ hopping terms, whose detailed form differs in different nonsymmorphic space groups (listed in SM). $\widetilde{h_{0}}({\bf k+Q})$ is obtained from $h_{0}({\bf k+Q})$ by correcting $\tau_{x,y}$ under time-reversal symmetry. The $\epsilon_{0}$ (or $t_{0}$) term includes the even (or odd) intra-sublattice hopping terms under $\bf k\rightarrow k+Q$, such as $\cos k_{y}$ (or $\cos k_{x}$) in the pedagogical example. As a next-nearest neighbor hopping, $t_{0}$ term does not significantly affect the magnitude of the odd-parity spin splitting, and the subsequent analysis assumes $t_{0}=0$. The three AFM states are described by $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},\hat{x}),J(\hat{x},0),\text{and }J(\hat{x},\hat{y})$. We note that this Hamiltonian can always be block-diagonalized since $[\rho_{z}\sigma_{z},H]=0$, allowing for analytical results.
Non-relativistic odd-parity spin-splittings: The odd-parity spin splitting can be obtained from the dispersion:
$\begin{split}&E_{\alpha\beta\gamma}=\epsilon_{0}+\gamma\bigl{\{}\vec{J}_{1}^{2}+\vec{J}_{2}^{2}+|{\bf t}|^{2}+2\beta\bigl{[}(t_{1}t_{3}+t_{2}t_{4})^{2}\\ &+(\vec{J}_{1}^{2}+\vec{J}_{2}^{2})|{\bf t}|^{2}/2+(\vec{J}_{1}^{2}-\vec{J}_{2}^{2})(t_{1}^{2}+t_{2}^{2}-t_{3}^{2}-t_{4}^{2})/2\\ &+(\vec{J}_{1}\cdot\vec{J}_{2})^{2}+2\alpha(t_{1}t_{4}-t_{2}t_{3})|\vec{J}_{1}\times\vec{J}_{2}|\bigr{]}^{1/2}\bigr{\}}^{1/2},\end{split}$ (4)
where $\alpha,\beta,\gamma=\pm$ and $|{\bf t}|^{2}\equiv t_{1}^{2}+t_{2}^{2}+t_{3}^{2}+t_{4}^{2}$. The spin splitting is between $\alpha=\pm$ bands. For the first two AFM states, $E_{\alpha\beta\gamma}$ is independent of $\alpha$, implying doubly degenerate bands. For the coplanar state, the magnitude of spin splitting depends on $J\sqrt{t_{1}t_{4}-t_{2}t_{3}}\propto J(t_{x}\widetilde{t_{y}}-\widetilde{t_{x}}t_{y})^{1/2}$. These spin splittings for different space groups and AFM wavevectors are shown in SM.
We now consider SG129(P4/nmm) with Wyckoff position 2a, relevant for FeSe (and 2c, for CeNiAsO Birk Hellenes et al. (2023)). The coefficients in the Hamiltonian are: $t_{x}=t_{x0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$ and $t_{y}=t_{y0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}(\cos k_{x}-\cos k_{y})\sin k_{z}$ (for 2c, remove $\cos k_{x}-\cos k_{y}$ factors here and below). For the AFM state with ${\bf Q}=(\pi,0,0)$, the other coefficients are $\widetilde{t_{x}}=-t_{y0}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}(\cos k_{x}-\cos k_{y})\sin k_{z}$ and $\widetilde{t_{y}}=-t_{x0}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$, from which we get $p$-wave spin splitting proportional to $\sin k_{x}$. For the AFM state at $(\pi,\pi,0)$, the other coefficients are $\widetilde{t_{x}}=t_{x0}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}$ and $\widetilde{t_{y}}=-t_{y0}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}(\cos k_{x}-\cos k_{y})\sin k_{z}$, from which we obtain an $h$-wave spin splitting proportional to $\sin k_{x}\sin k_{y}(\cos k_{x}-\cos k_{y})\sin k_{z}$ found in Fig. 2.
Non-relativistic Edelstein Effect: The Edelstein effect refers to the accumulation of spin-density induced by electric fields/currents, and serves as an important mechanism for charge-to-spin conversion. Traditionally, the Edelstein effect is relevant for systems with strong SOC, but has recently been proposed as a characteristic feature of certain non-collinear magnetic structures. This heralds the concept of a non-relativistic Edelstein effect González-Hernández et al. (2024); Hu et al. (2024); Chakraborty et al. (2024). Time-reversal symmetric odd-parity magnets are prime examples featuring the non-relativistic Edelstein effect. To demonstrate this for a specific case, we focus on the $p$-wave material candidate CeNiAsO Birk Hellenes et al. (2023); Hu et al. (2024). In Fig. 4(a,b) we show the minimal band structure and the associated normal state Fermi surface obtained from tight-binding fits to the low-energy bands from DFT calculations without SOC, see SM. The spin-polarization of the bands in the $p$-wave magnetic state is shown in Fig. 4(d).
We compute the Edelstein response tensor using the Boltzmann equation in the relaxation-time approximation, going to linear order in the electric field. This gives the transport coefficient $\chi_{ij}=\chi_{0}\sum_{\nu}\int d^{2}k\ (-\partial_{\epsilon}f_{0})\langle\sigma_{i\nu}\rangle(k)v_{j}$, where $\chi_{0}=\frac{\tau|e|\mu_{B}\mathcal{A}}{4\pi^{2}}$, $\nu$ is the band index, $v_{i}=\partial_{k_{i}}\epsilon(k)$, $f_{0}$ is the Fermi-Dirac distribution function and $\langle\sigma_{j\nu}\rangle(k)$ is the expectation value of the Pauli spin matrix in direction $j$. For $\chi_{0}$, $\tau$ denotes the relaxation time, $\mu_{B}$ is the Bohr magneton and $\mathcal{A}=a_{x}a_{y}$ is the area of the unit cell. Calculating this response for the bands shown in Fig. 4(c) gives $\chi_{zx}$ displayed in Fig. 4(c). As seen, most of the non-relativistic Edelstein response originates from the first band (black line), and we observe that the response changes sign as a function of the chemical potential Hu et al. (2024).
Figure 4: (a) Band structure relevant for CeNiAsO with two pairs of single-orbital models (black/gray). (b) Corresponding Fermi surface cut at $k_{z}=0$ in the BZ of the nonmagnetic system. (c) Edelstein susceptibility $\chi_{zx}$ as a function of chemical potential $\mu$ for different $p$-wave order parameters $J$, displaying the contributions of each band. (d) Fermi surface with the expectation value of the spin operator in the magnetic state exhibiting odd-parity spin splitting.
Nematic Order: Our microscopic model also reveals nematic order in the electronic spectrum. This appears in Eq. (4) as the coefficient of $|\vec{J}_{1}|^{2}-|\vec{J}_{2}|^{2}$, which is $(t_{1}^{2}+t_{2}^{2})-(t_{3}^{2}+t_{4}^{2})=t_{x}\widetilde{t_{x}}+t_{y}\widetilde{t_{y}}$. Specifically, we consider the well-known example of Fe-based superconductors with SG129 with ${\bf Q}=(\pi,\pi,0)$. Here $t_{x}=t_{x0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$ so that $\widetilde{t_{x}}=t_{x0}\sin\frac{k_{x}}{2}\sin\frac{k_{y}}{2}$. This yields a $\sqrt{Jt_{x0}}\sin k_{x}\sin k_{y}$-like nematic term in the dispersion. In the SM, we identify 22 materials in the Magndata database that should exhibit a similar nematic-driven electronic anisotropy.
Spin-scalar odd-parity state: We now consider the scalar odd-parity state. We will focus on state 1 in class 1 below, and the analysis for state 2 in class 2 is provided in SM. Although the state $(\vec{J}_{1},\vec{J}_{2})=J(\hat{x},\hat{x})$ has no spin-splitting [see Eq. (4)] and does not exhibit any symmetry breaking in the electronic dispersion, it can generate a non-relativistic Berry curvature dipole (unlike the spin-vector or nematic states), strongly contributing to nonlinear I-V transport. The Berry curvature is:
$\begin{split}\Omega^{ij}_{\alpha\beta\gamma}&=\frac{\beta}{|\vec{h}|^{3}}\sum_{abc}\epsilon_{abc}[h_{a}\partial_{k_{i}}h_{b}\partial_{k_{j}}h_{c}+\widetilde{h}_{a}\partial_{k_{i}}\widetilde{h}_{b}\partial_{k_{j}}\widetilde{h}_{c}],\\ \vec{h}&\equiv(Jt_{x},Jt_{y},J^{2}+(-t_{x}^{2}-t_{y}^{2}+\widetilde{t}_{x}^{2}+\widetilde{t}_{y}^{2})/4),\\ \vec{\widetilde{h}}&\equiv(J\widetilde{t}_{x},J\widetilde{t}_{y},J^{2}+(+t_{x}^{2}+t_{y}^{2}-\widetilde{t}_{x}^{2}-\widetilde{t}_{y}^{2})/4).\end{split}$ (5)
The derivation is in the SM. $\epsilon_{abc}$ is the antisymmetric tensor. The nontrivial k-odd dependence from $t_{x}t_{y}$ carries the site symmetry, as the AFM state distinguishes the two sublattices (state 1 in Fig. 3). This k-odd Berry curvature can generate non-zero Berry curvature dipoles $\langle d\Omega^{ij}/dk_{m}\rangle_{FS}$, which drives non-relativistic nonlinear I-V transport coefficients Sodemann and Fu (2015); Kaplan et al. (2024). In addition, this state is related to multiferroicity Cheong and Mostovoy (2007); Matsuda et al. (2024) and will also exhibit ferroelectric-like orderCheong and Mostovoy (2007).
Conclusions: We have developed a group-theory-based framework for generating odd-parity, time-reversal-preserving spin splitting from coplanar AFM states without SOC. By analyzing phenomenological models of 421 period-doubling AFM systems in nonsymmorphic space groups, we constructed minimal microscopic models for 119 of them, identifying three competing ground states: odd-parity spin splitting, nematic order, and scalar odd-parity order. Our DFT calculations demonstrate robust $h$-wave spin splitting in FeSe, characterized by a significant spin-splitting energy. Additionally, our minimal microscopic model reveals the non-relativistic Edelstein response in odd-parity spin-splitting states and non-relativistic nonlinear transport in scalar odd-parity states. We focused on applications to CeNiAsO and FeSe, with the model applicable to 67 AFM materials in Magndata database, providing a broad framework for spintronics application in realistic systems.
Acknowledgments We thank Philip Brydon, Rafael Fernandes, Ian Fisher, and Changhee Lee for useful discussions. M. R. acknowledges support from the Novo Nordisk Foundation grant NNF20OC0060019. A. K. acknowledges support by the Danish National Committee for Research Infrastructure (NUFI) through the ESS-Lighthouse Q-MAT. D. F. A. and Y. Y. were supported by the U.S. Department of Energy, Office of Basic Energy Sciences, Division of Materials Sciences and Engineering under Award No. DE-SC0021971 for symmetry-based calculations. M. W. and T. S. were supported by National Science Foundation Grant No. DMREF 2323857 for DFT calculations.
Appendix A Supplementary Material
Appendix B List of tables
Table.1 contains the symmetry of the odd-parity spin splitting, enforced by the nonsymmorphic space group symmetries.
Table.2 contains the symmetry of the nematicity, enforced by the nonsymmorphic space group symmetries.
Table.3 contains the symmetry of the scalar odd-parity order, enforced by the nonsymmorphic space group symmetries.
Table.4 contains the microscopic tight-binding parameters and the resulting spin splittings, for the class 1, where two atoms are related by inversion.
Table.5 contains the microscopic tight-binding parameters and the resulting spin splittings, for the class 2, where two atoms are at inversion center.
Appendix C Class 1
Here, we provide the detailed step to get the dispersion and Berry curvature for class 1, where the atoms related by inversion. The Hamiltonian is:
$\begin{split}H=t_{1}\rho_{0}\tau_{x}+t_{2}\rho_{0}\tau_{y}+t_{3}\rho_{z}\tau_{x}+t_{4}\rho_{z}\tau_{y}+\rho_{x}\vec{J_{1}}\cdot\vec{\sigma}+\rho_{x}\tau_{z}\vec{J_{2}}\cdot\vec{\sigma},\end{split}$ (6)
where $t_{1}\equiv(t_{x}+\widetilde{t_{x}})/2$, $t_{2}\equiv(t_{y}+\widetilde{t_{y}})/2$, $t_{3}\equiv(t_{x}-\widetilde{t_{x}})/2$, and $t_{4}\equiv(t_{y}-\widetilde{t_{y}})/2$. We have taken $t_{0}=0$ and dropped the trivial $\epsilon_{0}$ term for simplicity.
The three AFM states are described by $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},\hat{x}),J(\hat{x},0),\text{and }J(\hat{x},\hat{y})$. Without loss of generality, we will consider $\vec{J}_{1,2}$ in the xy plane. Using $[\rho_{z}\sigma_{z},H]=0$, we can block-diagonalize the Hamiltonian:
$\begin{split}&u^{\dagger}Hu=\left(\begin{array}[]{cc}H_{\alpha=+}&\\ &H_{\alpha=-}\end{array}\right),\;u=\left(\begin{array}[]{cccc}1&&&\\ &&1&\\ &&&1\\ &1&&\end{array}\right),\\ &H_{\alpha}=t_{1}\widetilde{\rho_{0}}\tau_{x}+t_{2}\widetilde{\rho_{0}}\tau_{y}+t_{3}\widetilde{\rho_{z}}\tau_{x}+t_{4}\widetilde{\rho_{z}}\tau_{y}+\vec{J}_{1\alpha}\cdot\vec{\widetilde{\rho}}+\tau_{z}\vec{J}_{2\alpha}\cdot\vec{\widetilde{\rho}},\end{split}$ (7)
where the unitary transformation $u$ is in $(\rho,\sigma)$ space, and $\vec{J}_{i\alpha}\equiv(J_{ix},\alpha J_{iy},0)$. Using $\{\widetilde{\rho_{z}}\tau_{z},H_{\alpha}\}=0$, we can further off-block-diagonalize the resulting $4\times 4$ Hamiltonian:
$\begin{split}\widetilde{H}_{\alpha}&=\widetilde{u}^{\dagger}H_{\alpha}\widetilde{u}=\left(\begin{array}[]{cc}0&M\\ M^{\dagger}&0\end{array}\right),\;\widetilde{u}=\left(\begin{array}[]{cccc}1&&&\\ &&1&\\ &&&1\\ &1&&\end{array}\right),\\ M&\equiv\left(\begin{array}[]{cc}t_{1}-it_{2}+t_{3}-it_{4}&J_{1x}+J_{2x}-i\alpha J_{1y}-i\alpha J_{2y}\\ J_{1x}-J_{2x}+i\alpha J_{1y}-i\alpha J_{2y}&t_{1}+it_{2}-t_{3}-it_{4}\end{array}\right),\end{split}$ (8)
The eigenvectors $[v_{1},v_{2}]^{T}$ of $\widetilde{H}_{\alpha}$ satisfy:
$\begin{split}M^{\dagger}Mv_{2}=E^{2}v_{2},\;v_{1}=\frac{1}{E}Mv_{2}.\end{split}$ (9)
Notably, $v_{1}$ also satisfies $MM^{\dagger}v_{1}=E^{2}v_{1}$. Here,
$\begin{split}M^{\dagger}M&=|\vec{J}_{1}|^{2}+|\vec{J}_{2}|^{2}+t_{1}^{2}+t_{2}^{2}+t_{3}^{2}+t_{4}^{2}+2(t_{1}J_{1x}+\alpha t_{2}J_{1y}+t_{3}J_{2x}+\alpha t_{4}J_{2y})\widetilde{\tau_{1}}\\ &+2(\alpha t_{1}J_{1y}-t_{2}J_{1x}+\alpha t_{3}J_{2y}-t_{4}J_{2x})\widetilde{\tau_{2}}+2(-\vec{J}_{1}\cdot\vec{J}_{2}+t_{1}t_{3}+t_{2}t_{4})\widetilde{\tau_{3}}\\ MM^{\dagger}&=|\vec{J}_{1}|^{2}+|\vec{J}_{2}|^{2}+t_{1}^{2}+t_{2}^{2}+t_{3}^{2}+t_{4}^{2}+2(t_{1}J_{1x}-\alpha t_{2}J_{1y}-t_{3}J_{2x}+\alpha t_{4}J_{2y})\widetilde{\tau_{1}}\\ &+2(\alpha t_{1}J_{1y}+t_{2}J_{1x}-\alpha t_{3}J_{2y}-t_{4}J_{2x})\widetilde{\tau_{2}}+2(+\vec{J}_{1}\cdot\vec{J}_{2}+t_{1}t_{3}+t_{2}t_{4})\widetilde{\tau_{3}}.\end{split}$ (10)
One can then reproduce the dispersion by diagonalizing $M^{\dagger}M$:
$\begin{split}E_{\alpha\beta\gamma}&=\epsilon_{0}+\gamma\bigl{\{}\vec{J}_{1}^{2}+\vec{J}_{2}^{2}+|{\bf t}|^{2}+2\beta\bigl{[}(t_{1}t_{3}+t_{2}t_{4})^{2}+(\vec{J}_{1}^{2}+\vec{J}_{2}^{2})|{\bf t}|^{2}/2\\ &+(\vec{J}_{1}^{2}-\vec{J}_{2}^{2})(t_{1}^{2}+t_{2}^{2}-t_{3}^{2}-t_{4}^{2})/2+(\vec{J}_{1}\cdot\vec{J}_{2})^{2}+2\alpha(t_{1}t_{4}-t_{2}t_{3})|\vec{J}_{1}\times\vec{J}_{2}|\bigr{]}^{1/2}\bigr{\}}^{1/2},\end{split}$ (11)
We now turn to the Berry curvature. Notably, $v_{1}$ and $v_{2}$ have the same norm:
$\begin{split}v_{1}^{\dagger}v_{1}=\frac{1}{E^{2}}v_{2}^{\dagger}M^{\dagger}Mv_{2}=\frac{1}{E^{2}}v_{2}^{\dagger}(E^{2})v_{2}=v_{2}^{\dagger}v_{2}.\end{split}$ (12)
That says, after normalizing the eigenvector, $v_{1}^{\dagger}v_{1}=v_{2}^{\dagger}v_{2}=1/2$. The normalization of $v_{1}$ and $v_{2}$ are thus independent. The Berry connection (and Berry curvature) can be decomposed into contributions from $v_{1}$ and $v_{2}$:
$\begin{split}\vec{A}_{i}=i[v_{1}^{\dagger},v_{2}^{\dagger}]\partial_{i}[v_{1},v_{2}]=iv_{1}^{\dagger}\partial_{i}v_{1}+iv_{2}^{\dagger}\partial_{i}v_{2}.\end{split}$ (13)
The contribution from $v_{1}$ (or $v_{2}$) can be determined from $MM^{\dagger}$ (or $M^{\dagger}M$). For a general $2\times 2$ Hamiltonian $H=h_{1}\tau_{x}+h_{2}\tau_{y}+h_{3}\tau_{z}$, the Berry curvature is $\Omega_{ij}=\frac{1}{2h^{3}}\sum_{a,b,c=1}^{3}\epsilon_{abc}h_{a}\partial_{i}h_{b}\partial_{j}h_{c}$. We can then apply this formula to obtain the contribution from $v_{1}$ and $v_{2}$, with an extra prefactor of $1/4$ from their normalization. For the nematic state $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},0)$ and odd-parity spin-splitting state $(\vec{J}_{1},\vec{J}_{2})=J(\hat{x},\hat{y})$, these two contribution exactly cancel each other, leading to zero Berry curvature. For the scalar odd-parity state $(\vec{J}_{1},\vec{J}_{2})=J(\hat{x},\hat{x})$, the Berry curvature is
$\begin{split}\Omega^{ij}_{\alpha\beta\gamma}=\frac{\beta}{|\vec{h}|^{3}}\sum_{abc}\epsilon_{abc}[h_{a}\partial_{k_{i}}&h_{b}\partial_{k_{j}}h_{c}+\widetilde{h}_{a}\partial_{k_{i}}\widetilde{h}_{b}\partial_{k_{j}}\widetilde{h}_{c}]\\ \vec{h}\equiv(Jt_{x},Jt_{y},J^{2}+(-t_{x}^{2}-t_{y}^{2}+\widetilde{t}_{x}^{2}+\widetilde{t}_{y}^{2})/4),&\;\;\vec{\widetilde{h}}\equiv(J\widetilde{t}_{x},J\widetilde{t}_{y},J^{2}+(+t_{x}^{2}+t_{y}^{2}-\widetilde{t}_{x}^{2}-\widetilde{t}_{y}^{2})/4)\end{split}$ (14)
The inversion symmetry breaking is on the order of $J^{2}$, depending on $t_{x}t_{y}-\widetilde{t_{x}}\widetilde{t_{y}}$. For example, in SG129(2c), $t_{x}=t_{x0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$ and $t_{y}=t_{y0}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\sin k_{z}$. With ${\bf Q}=(\pi,0,0)$, $\widetilde{t_{x}}=-t_{x0}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}\sin k_{z}$ and $\widetilde{t_{y}}=-t_{u0}\sin\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$. We obtain $(t_{x}t_{y}-\widetilde{t_{x}}\widetilde{t_{y}})\propto\sin k_{z}$, in agreement with the symmetry requirements from the phenomenological result in Table.3.
Appendix D Class 2
In this section, we repeat the above analysis for the second inversion class, where the atoms are located at inversion center. The Hamiltonian of the doubled unit cell is:
$\begin{split}H&=\begin{pmatrix}h_{0}({\bf k})&O_{M}\\ O_{M}^{\dagger}&\widetilde{h_{0}}({\bf k+Q})\end{pmatrix}=\begin{pmatrix}\epsilon_{0}+t_{0}+t_{x}\tau_{x}+t_{z}\tau_{z}&\vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma}\\ \vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma}&\epsilon_{0}-t_{0}+\widetilde{t}_{y}\tau_{y}+\widetilde{t}_{z}\tau_{z}\end{pmatrix}\\ &\equiv\epsilon_{0}+t_{1}\rho_{0}\tau_{x}+t_{2}\rho_{0}\tau_{y}+t_{3}\rho_{z}\tau_{x}+t_{4}\rho_{z}\tau_{y}+t_{5}\rho_{0}\tau_{z}+t_{6}\rho_{z}\tau_{z}+t_{0}\rho_{z}+\rho_{x}(\vec{J}_{1}\cdot\vec{\sigma}+\tau_{z}\vec{J}_{2}\cdot\vec{\sigma})\end{split}$ (15)
where $t_{1}=t_{3}\equiv t_{x}/2$, $t_{2}=-t_{4}\equiv\widetilde{t_{y}}/2$, $t_{5}\equiv(t_{z}+\widetilde{t_{z}})/2$, and $t_{6}\equiv(t_{z}-\widetilde{t_{z}})/2$. Pauli matrix $\tau_{z}=\pm$ distinguishes sublattices within each nonmagnetic unit cell, and $\rho_{x}=\pm$ differentiates nonmagnetic unit cells when unit cell doubling is applied. Time-reversal operator and inversion operator are $T=i\sigma_{y}K$ and $I=\frac{1+\rho_{z}}{2}\otimes\tau_{0}+\frac{1-\rho_{z}}{2}\otimes\tau_{z}$, with momentum flip.
$t_{x}$ and $t_{z}$ are k-even hopping terms in the nonmagnetic unit cells, listed in Table.5. $t_{x}$ is the nearest neighbored hopping, and $t_{z}$ hopping is generically weaker, since it comes from further neighbored hopping. For every k-even $t_{x}\tau_{x}$ terms in $h_{0}({\bf k})$, we obtain the k-odd $\widetilde{t_{y}}\tau_{y}$ term from $h_{0}({\bf k+Q})$, with the correct operators $\tau_{y}$ due to time-reversal symmetry. For every k-even $t_{z}\tau_{z}$ term in $h_{0}({\bf k})$, we obtain the k-even $\widetilde{t_{z}}\tau_{z}$ term from $h_{0}({\bf k+Q})$. In the following analysis, we will take $t_{0}=0$ and drop the trivial $\epsilon_{0}$ term for simplicity.
The three AFM states are described by $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},\hat{x}),J(\hat{x},0),\text{and }J(\hat{x},\hat{y})$.Without loss of generality, we will consider $\vec{J}_{1,2}$ in the xy plane. Using $[\rho_{z}\sigma_{z},H]=0$, we can block-diagonalize the Hamiltonian:
$\begin{split}&u^{\dagger}Hu=\left(\begin{array}[]{cc}H_{\alpha=+}&\\ &H_{\alpha=-}\end{array}\right),\;u=\left(\begin{array}[]{cccc}1&&&\\ &&1&\\ &&&1\\ &1&&\end{array}\right),\\ &H_{\alpha}=t_{1}\widetilde{\rho_{0}}\tau_{x}+t_{2}\widetilde{\rho_{0}}\tau_{y}+t_{3}\widetilde{\rho_{z}}\tau_{x}+t_{4}\widetilde{\rho_{z}}\tau_{y}+t_{5}\widetilde{\rho_{0}}\tau_{z}+t_{6}\widetilde{\rho_{z}}\tau_{z}+t_{0}\widetilde{\rho_{z}}+\vec{J}_{1\alpha}\cdot\vec{\widetilde{\rho}}+\tau_{z}\vec{J}_{2\alpha}\cdot\vec{\widetilde{\rho}},\end{split}$ (16)
where the unitary transformation $u$ is in $(\rho,\sigma)$ space, and $\vec{J}_{i\alpha}\equiv(J_{ix},\alpha J_{iy},0)$.
For the first two AFM orders, the band remains doubly degenerate. For the ${\bf S_{1}}\times{\bf S_{2}}\neq 0$ state, the dispersion is:
$\begin{split}E_{\alpha\beta\gamma}=\gamma\left[2J^{2}+|{\bf t}|^{2}+2\beta\sqrt{J^{2}(|{\bf t}|^{2}+t_{5}^{2}-t_{6}^{2})+2\alpha J^{2}(t_{1}t_{4}-t_{2}t_{3})+(t_{1}t_{3}+t_{2}t_{4}+t_{5}t_{6})^{2}}\right]^{1/2},\end{split}$ (17)
where $\alpha,\beta,\gamma=\pm$ and $|{\bf t}|^{2}\equiv t_{1}^{2}+t_{2}^{2}+t_{3}^{2}+t_{4}^{2}+t_{5}^{2}+t_{6}^{2}$. The magnitude of spin splitting depends on $J\sqrt{t_{1}t_{4}-t_{2}t_{3}}\propto J(t_{x}\widetilde{t_{y}})^{1/2}$.
We now turn to the Berry curvature dipole of the collinear scalar odd-parity state $(\vec{J}_{1},\vec{J}_{2})=(\frac{{\bf S_{1}+S_{2}}}{2},\frac{{\bf S_{1}-S_{2}}}{2})=J(\hat{x},0)$, the dispersion is
$\begin{split}E_{\beta\gamma}=\gamma\left[J^{2}+|{\bf t}|^{2}+2\beta\sqrt{J^{2}(t_{1}^{2}+t_{2}^{2}+t_{5}^{2})+(t_{1}t_{3}+t_{2}t_{4}+t_{5}t_{6})^{2}}\right]^{1/2},\end{split}$ (18)
[…]
$\begin{split}&P_{\alpha\beta\gamma}=\frac{1}{4}\left[1+\frac{H_{\alpha}}{E_{\beta\gamma}}\right]\left[1+\frac{\bar{H}_{\alpha}}{\bar{E}_{\beta}}\right],\\ &\bar{H}_{\alpha}\equiv 2\left[(t_{1}t_{3}+t_{2}t_{4}+t_{5}t_{6})\widetilde{\rho_{z}}\tau_{0}+Jt_{1}\widetilde{\rho_{x}}\tau_{x}+Jt_{2}\widetilde{\rho_{x}}\tau_{y}+Jt_{5}\widetilde{\rho_{x}}\tau_{z}\right],\\ &\bar{E}_{\beta}\equiv 2\beta\sqrt{J^{2}(t_{1}^{2}+t_{2}^{2}+t_{5}^{2})+(t_{1}t_{3}+t_{2}t_{4}+t_{5}t_{6})^{2}}\end{split}$ (19)
The Berry curvature is
$\begin{split}&\bar{\beta}=+:\Omega^{ij}_{\alpha\beta\gamma}=\frac{J^{2}}{16\bar{E}_{\beta}^{3}E^{3}_{\alpha\beta\gamma}}[\partial_{i}t_{z}(3t_{x}^{5}\partial_{j}\widetilde{t}_{y}-9t_{x}^{4}\widetilde{t}_{y}\partial_{j}t_{x}+6t_{x}^{3}\widetilde{t}_{y}^{2}\partial_{j}\widetilde{t}_{y}+2t_{x}^{2}\widetilde{t}_{y}^{3}\partial_{j}t_{x}-t_{x}\widetilde{t}_{y}^{4}\partial_{j}\widetilde{t}_{y}-\widetilde{t}_{y}^{5}\partial_{j}t_{x})\\ &-t_{z}\partial_{i}t_{x}\partial_{j}\widetilde{t}_{y}(3t_{x}^{4}+6t_{x}^{2}\widetilde{t}_{y}^{2}-\widetilde{t}_{y}^{4})-8\widetilde{t_{z}}t_{x}^{4}\partial_{i}t_{x}\partial_{j}\widetilde{t}_{y}+\partial_{i}\widetilde{t_{z}}(4t_{x}^{5}\partial_{j}\widetilde{t}_{y}-8t_{x}^{4}\widetilde{t}_{y}\partial_{j}t_{x}+4t_{x}^{3}\widetilde{t}_{y}^{2}\partial_{j}\widetilde{t}_{y})-(i\leftrightarrow j)],\\ &\bar{\beta}=-:\Omega^{ij}_{\alpha\beta\gamma}=\frac{J^{2}}{16\bar{E}_{\beta}^{3}E^{3}_{\alpha\beta\gamma}}[\partial_{i}\widetilde{t_{z}}(-3\widetilde{t}_{y}^{5}\partial_{j}t_{x}+9t_{x}\widetilde{t}_{y}^{4}\partial_{j}\widetilde{t}_{y}-6t_{x}^{2}\widetilde{t}_{y}^{3}\partial_{j}t_{x}-2t_{x}^{3}\widetilde{t}_{y}^{2}\partial_{j}\widetilde{t}_{y}+t_{x}^{4}\widetilde{t}_{y}\partial_{j}t_{x}+t_{x}^{5}\partial_{j}\widetilde{t}_{y})\\ &-\widetilde{t_{z}}\partial_{i}t_{x}\partial_{j}\widetilde{t}_{y}(3\widetilde{t}_{y}^{4}+6t_{x}^{2}\widetilde{t}_{y}^{2}-t_{x}^{4})-8t_{z}\widetilde{t}_{y}^{4}\partial_{i}t_{x}\partial_{j}\widetilde{t}_{y}+\partial_{i}t_{z}(-4\partial_{j}t_{x}\widetilde{t}_{y}^{5}+8\partial_{j}\widetilde{t}_{y}t_{x}\widetilde{t}_{y}^{4}-4\partial_{j}t_{x}t_{x}^{2}\widetilde{t}_{y}^{3})-(i\leftrightarrow j)]\end{split}$ (20)
where $\bar{\beta}\equiv\beta\text{sgn}(t_{x}^{2}-\widetilde{t}_{y}^{2})$. We have kept the leading contributions, quadratic in $J$ and linear in the further hopping strength $t_{z}$ (or $\widetilde{t_{z}}$). The symmetry of Berry curvature follows $J^{2}t_{x}\widetilde{t_{y}}t_{z}$ and $J^{2}t_{x}\widetilde{t_{y}}\widetilde{t_{z}}$. Such k-odd Berry curvature allows non-zero Berry curvature dipoles $\langle d\Omega^{ij}/dk_{m}\rangle_{FS}$.
Appendix E List of materials
In this section, we list the materials where the Landau theory is applicable. The symbol * denotes materials where certain magnetic atoms can be described by the microscopic models. Space groups, wavevectors, and Magndata entry labels are included. The category Scalar and/or spin odd-parity order appears because SOC makes the precise categorization of the non-relativistic state uncertain and both possibilities are consistent with the experimental structure.
Spin odd-parity DyMn2O5(55X,1.324), Tm5Pt2In4(55T,1.790*), CoNb2O6(60Y,1.224),
La3/8Ca5/8MnO3(62Z,1.173), La1/3Ca2/3MnO3(62Z,1.174),
La1/3Ca2/3MnO3(62X,1.175), Gd2BaCuO5(62Z,1.443), Er2Pt(62X,1.444),
CeNiAsO(129X,1.272*), Sr2FeO3Cl(129M,1.380,1.382*), Sr2FeO3Br(129M,1.381,1.383*),
Sr2FeO3F(129M,1.385,1.386,1.387*), NiCr2O4(141M,1.688), Dy2Co3Al9(63Z,1.267)
Scalar odd-parity Na2MnF5(14Z,1.55*), Lu2MnCoO6(14A,1.32*), YbLuCoMnO6(14B,1.329*),
Lu2CoMnO6(14B,1.330*), Yb2CoMnO6(14B,1.328*), SmMn2O5(55X,1.192),
DyMn2O5 (55X,1.599,1.76), PrMn2O5(55X,1.325), Sr2Fe3Se2O3(55U,1.463,1.626),
Tb5Pd2In4(55Y,1.697*), FeNb2O6(60Y,1.655), BaCdVO(PO4)2(61Y,1.298),
LuMnO3(62X,1.101), HoMnO3(62X,1.20), PrNiO3(62T,1.43), NdNiO3(62T,1.45),
TmMnO3(62X,1.341), SmNiO3(62T,1.353), EuNiO3(62T,1.354), HoNiGe(62T,1.374),
Na2CuSO4Cl2(62X,1.682), NiCr2O4(141M,1.685), Na2Ni2TeO6(193L,1.646*), MnS2(205X,1.18)
Scalar and/or spin odd-parity PrMn2O5(55X,1.19), GdMn2O5(55X,1.54,1.299,1.300), BiMn2O5(55U,1.74,1.75),
Tm5Ni2In4(55T,1.170*), NdNiO3(62T,1.44), Cs2CoCl4(62T,1.51),LuMnO3(62X,1.340)
Nematic Fe(ND3)2PO4 (14B,1.66), NaMnGe2O6(15A,1.260), PrFe2Al8(55U,1.681), VOCl(59R,1.37),
MnV2O6(60Z,1.196), NiNb2O6(60Y,1.654), CoNb2O6(60Z,1.656), BaNd2O4(62T,1.95),
ErNiGe(62Y,1.379), Y2BaCuO5(62T,1.445), SrNd2O4(62T,1.577), DyBaCuO5(62Y,1.650),
Dy2TiO5(62Y,1.698), PrPdSn(62T,1.744) , Ho2BaCuO5(62Y,1.651), DyGe(63Y,1.361)
Li2VOSiO4(129M,1.9), Sr2CoO3Cl(129M,1.389), Fe1.05Te(129R,1.434),
[…]
Table 1: Little co-group, momentum-dependent spin-splitting, space group and wavevectors. Labels denote spin splittings not along the conventional axis. For instance, $11C_{1}$ has spin splitting along $k_{y}$ instead of $k_{z}$.
$P$ Spin-splitting Space group & Wavevector
$C_{2h}$ $k_{z}$ $11(C^{*}_{1},D^{*}_{1},E^{*}_{1},Z^{*}_{1}),14(C^{*}_{1},Z^{*}_{1}),63(R_{1}),176(L_{1})$
$C_{4h}$ $k_{z}$ $84(A_{1},Z_{1}),85(A_{1},M_{1}),86(M_{1},Z_{1}),88(M_{1})$
$C_{6h}$ $k_{z}$ $176(A_{1})$
$D_{2h}$ $k_{z,y,x}$ $\begin{matrix}48(S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),49(R_{1,2},T_{1,2},U_{1,2},Z_{1,2}),50(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),\\ 51(R_{1,2},S_{1,2},U_{1,2},X_{1,2}),52(R_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),53(S_{1,2},T_{1,2},X_{1,2},Z_{1,2}),\\ 54(S_{1,2},T_{1,2},X_{1,2},Z_{1,2}),55(T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),56(X_{1,2},Y_{1,2},Z_{1,2}),57(S_{1,2},U_{1,2},Y_{1,2},Z_{1,2}),\\ 58(X_{1,2},Y_{1,2},Z_{1,2}),59(T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),60(X_{1,2},Y_{1,2},Z_{1,2}),61(X_{1,2},Y_{1,2},Z_{1,2}),\\ 62(X_{1,2},Y_{1,2},Z_{1,2}),63(T_{1,2},Z_{1,2}),64(T_{1,2},Z_{1,2}),66(T_{1,2},Z_{1,2}),68(T_{1,2},Z_{1,2}),70(T_{1,2},Y_{1,2},Z_{1,2}),\\ 124(R_{1,2}),125(R_{1,2},X_{1,2}),126(R_{1,2},X_{1,2}),127(R_{1,2},X_{1,2}),128(X_{1,2}),129(R_{1,2},X_{1,2}),130(X_{1,2}),\\ 132(R_{1,2}),133(R_{1,2},X_{1,2}),134(R_{1,2},X_{1,2}),135(R_{1,2},X_{1,2}),136(X_{1,2}),137(R_{1,2},X_{1,2}),138(X_{1,2}),\\ 141(X_{1,2}),142(X_{1,2}),192(L_{1,2}),193(L_{1,2}),194(L_{1,2}),201(M_{1,2},X_{1,2}),203(X_{1,2}),205(X_{1,2}),230(N_{1,2})\end{matrix}$
$D_{4h}$ $k_{z}$ $\begin{matrix}124(A_{1,2},Z_{1,2}),125(A_{3,4},M_{3,4}),126(A_{3,4},M_{3,4},Z_{1,2}),128(Z_{1,2}),130(Z_{1,2}),131(A_{1,2},Z_{1,2}),\\ 132(A_{1,2},Z_{1,2}),133(A_{1,2},M_{1,2},Z_{1,2}),134(M_{1,2},Z_{1,2}),135(Z_{1,2}),136(Z_{1,2}),137(Z_{1,2}),138(Z_{1,2}),\\ 141(M_{1,2}),142(M_{1,2}),222(M_{3,4},X^{*}_{1,2}),223(X^{*}_{1,2}),224(M_{1,2},X^{*}_{1,2}),227(X^{*}_{1,2}),228(X^{*}_{1,2})\end{matrix}$
$D_{3d}$ $k_{z}$ $163(A_{3}),165(A_{3}),167(T_{3}),226(L^{*}_{3}),228(L^{*}_{3})$
$D_{6h}$ $k_{z}$ $192(A_{5,6}),193(A_{1,2}),194(A_{1,2})$
$C_{2h}$ $c_{1}k_{x}+c_{2}k_{y}$ $\begin{matrix}13(A^{*}_{1},B^{*}_{1},D^{*}_{1},E^{*}_{1}),14(A^{*}_{1},B^{*}_{1}),15(A^{*}_{1},M^{*}_{1}),64(S_{1}),67(R_{1},S_{1}),68(R_{1},S_{1}),72(S^{*}_{1},R^{*}_{1}),73(T_{1},S^{*}_{1},R^{*}_{1}),\\ 74(T_{1}),85(R_{1},X_{1}),86(R_{1},X_{1}),88(X_{1}),140(N^{*}_{1}),142(N^{*}_{1}),163(L^{*}_{1}),165(L^{*}_{1}),167(L^{*}_{1}),206(N_{1})\end{matrix}$
$C_{4h}$ $\begin{matrix}c_{1}k_{x}k_{y}k_{z}+\\ c_{2}k_{z}(k_{x}^{2}-k_{y}^{2})\end{matrix}$ $84(A_{2},Z_{2}),85(A_{2},M_{2}),86(M_{2},Z_{2}),88(M_{2})$
$D_{4h}$ $k_{z}(k_{x}^{2}-k_{y}^{2})$ $\begin{matrix}125(A_{1,2},M_{1,2}),126(M_{1,2}),132(A_{3,4},Z_{3,4}),133(M_{3,4}),134(M_{3,4},Z_{3,4}),136(Z_{3,4}),\\ 138(Z_{3,4}),222(M_{1,2}),224(M_{3,4},X^{*}_{3,4}),227(X^{*}_{3,4}),228(X^{*}_{3,4})\end{matrix}$
$D_{2h}$ $k_{x}k_{y}k_{z}$ $56(S_{1,2}),58(R_{1,2}),59(R_{1,2},S_{1,2}),62(T_{1,2})$
$D_{4h}$ $k_{x}k_{y}k_{z}$ $\begin{matrix}126(A_{1,2}),128(A_{1,2}),129(A_{3,4},M_{3,4}),130(M_{3,4}),131(A_{3,4},Z_{3,4}),133(Z_{3,4}),135(Z_{3,4}),\\ 136(A_{1,2}),137(A_{1,2},M_{3,4},Z_{3,4}),138(M_{3,4}),141(M_{3,4}),142(M_{3,4}),223(X^{*}_{3,4})\end{matrix}$
$O_{h}$ $k_{x}k_{y}k_{z}$ $222(R_{1}),223(R_{1}),230(H_{1})$
$D_{4h}$ $\begin{matrix}k_{x}k_{y}k_{z}\times\\ (k_{x}^{2}-k_{y}^{2})\end{matrix}$ $129(A_{1,2},M_{1,2}),130(M_{1,2}),136(A_{3,4}),137(M_{1,2}),138(M_{1,2})$
[…]
$P$ IR of P Nematicity Space group & Wavevector
$C_{2h}$ $B_{g}$ $\begin{matrix}c_{1}k_{x}k_{z}\\ +c_{2}k_{y}k_{z}\end{matrix}$ $\begin{matrix}11(C^{*}_{1},D^{*}_{1},E^{*}_{1},Z^{*}_{1}),13(A^{*}_{1},B^{*}_{1},D^{*}_{1},E^{*}_{1}),14(A^{*}_{1},B^{*}_{1},C^{*}_{1},Z^{*}_{1}),15(A^{*}_{1},M^{*}_{1}),63(R_{1}),\\ 64(S_{1}),67(R_{1},S_{1}),68(R_{1},S_{1}),72(S^{*}_{1},R^{*}_{1}),73(T_{1},S^{*}_{1},R^{*}_{1}),74(T_{1}),85(R_{1},X_{1}),\\ 86(R_{1},X_{1}),88(X_{1}),140(N^{*}_{1}),142(N^{*}_{1}),163(L^{*}_{1}),165(L^{*}_{1}),167(L^{*}_{1}),176(L_{1}),206(N_{1})\end{matrix}$
$C_{4h}$ $B_{g}$ $\begin{matrix}c_{1}(k_{x}^{2}-k_{y}^{2})\\ +c_{2}k_{x}k_{y}\end{matrix}$ $84(A_{1,2},Z_{1,2}),85(A_{1,2},M_{1,2}),86(M_{1,2},Z_{1,2}),88(M_{1,2})$
$C_{6h}$ $B_{g}$ $k_{x}k_{z}(k_{x}^{2}-3k_{y}^{2})$ $176(A_{1})$
$D_{2h}$ $\begin{matrix}B_{1g}\\ B_{2g}\\ B_{3g}\end{matrix}$ $\begin{matrix}k_{x}k_{y}\\ k_{z}k_{x}\\ k_{y}k_{z}\end{matrix}$ $\begin{matrix}48(S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),49(R_{1,2},T_{1,2},U_{1,2},Z_{1,2}),\\ 50(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),51(R_{1,2},S_{1,2},U_{1,2},X_{1,2}),52(R_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),\\ 53(S_{1,2},T_{1,2},X_{1,2},Z_{1,2}),54(S_{1,2},T_{1,2},X_{1,2},Z_{1,2}),55(T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),\\ 56(S_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),57(S_{1,2},U_{1,2},Y_{1,2},Z_{1,2}),58(R_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),\\ 59(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),60(X_{1,2},Y_{1,2},Z_{1,2}),61(X_{1,2},Y_{1,2},Z_{1,2}),\\ 62(T_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),63(T_{1,2},Z_{1,2}),64(T_{1,2},Z_{1,2}),66(T_{1,2},Z_{1,2}),68(T_{1,2},Z_{1,2}),\\ 70(T_{1,2},Y_{1,2},Z_{1,2}),124(R_{1,2}),125(R_{1,2},X_{1,2}),126(R_{1,2},X_{1,2}),127(R_{1,2},X_{1,2}),128(X_{1,2}),\\ 129(R_{1,2},X_{1,2}),130(X_{1,2}),132(R_{1,2}),133(R_{1,2},X_{1,2}),134(R_{1,2},X_{1,2}),135(R_{1,2},X_{1,2}),\\ 136(X_{1,2}),137(R_{1,2},X_{1,2}),138(X_{1,2}),141(X_{1,2}),142(X_{1,2}),\\ 192(L_{1,2}),193(L_{1,2}),194(L_{1,2}),201(M_{1,2},X_{1,2}),203(X_{1,2}),205(X_{1,2}),230(N_{1,2})\end{matrix}$
$D_{3d}$ $A_{2g}$ $k_{x}k_{z}(k_{x}^{2}-3k_{y}^{2})$ $163(A_{3}),165(A_{3}),167(T_{3}),226(L^{*}_{3}),228(L^{*}_{3})$
$D_{4h}$ $B_{1g}$ $k_{x}^{2}-k_{y}^{2}$ $\begin{matrix}126(A_{1,2,3,4}),131(A_{1,2,3,4},Z_{1,2,3,4}),133(Z_{1,2,3,4}),135(Z_{1,2,3,4}),\\ 137(Z_{1,2,3,4}),141(M_{1,2,3,4}),142(M_{1,2,3,4}),223(X^{*}_{1,2,3,4})\end{matrix}$
$D_{4h}$ $B_{2g}$ $k_{x}k_{y}$ $\begin{matrix}125(A_{1,2,3,4},M_{1,2,3,4}),126(M_{1,2,3,4}),129(A_{1,2,3,4},M_{1,2,3,4}),130(M_{1,2,3,4}),\\ 132(A_{1,2,3,4},Z_{1,2,3,4}),133(M_{1,2,3,4}),134(M_{1,2,3,4},Z_{1,2,3,4}),136(A_{1,2,3,4},Z_{1,2,3,4}),\\ 137(M_{1,2,3,4}),138(M_{1,2,3,4},Z_{1,2,3,4}),222(M_{1,2,3,4}),\\ 224(M_{1,2,3,4},X^{*}_{1,2,3,4}),227(X^{*}_{1,2,3,4}),228(X^{*}_{1,2,3,4})\end{matrix}$
$D_{6h}$ $A_{2g}$ $\begin{matrix}k_{x}k_{y}(k_{x}^{2}-3k_{y}^{2})\\ \times(k_{y}^{2}-3k_{x}^{2})\end{matrix}$ $192(A_{5,6})$
$D_{6h}$ $B_{1g}$ $k_{y}k_{z}(k_{y}^{2}-3k_{x}^{2})$ $193(A_{1,2})$
$D_{6h}$ $B_{2g}$ $k_{x}k_{z}(k_{x}^{2}-3k_{y}^{2})$ $194(A_{1,2})$
$O_{h}$ $A_{2g}$ $\begin{matrix}(k_{x}^{2}-k_{y}^{2})(k_{y}^{2}-k_{z}^{2})\\ \times(k_{z}^{2}-k_{x}^{2})\end{matrix}$ $222(R_{1}),223(R_{1}),230(H_{1})$
Table 3: Little co-group, spin-scalar odd-parity state, space group and wavevectors.
$P$ IR of $P$ Symmetry Space group & Wavevector
$C_{2h}$ $A_{u}$ $z$ $\begin{matrix}13(A^{*}_{1},B^{*}_{1},D^{*}_{1},E^{*}_{1}),14(A^{*}_{1},B^{*}_{1}),15(A^{*}_{1},M^{*}_{1}),64(S_{1}),67(R_{1},S_{1}),68(R_{1},S_{1}),\\ 72(S^{*}_{1},R^{*}_{1}),73(T_{1},S^{*}_{1},R^{*}_{1}),74(T_{1}),85(R_{1},X_{1}),86(R_{1},X_{1}),88(X_{1}),140(N^{*}_{1}),\\ 142(N^{*}_{1}),163(L^{*}_{1}),165(L^{*}_{1}),167(L^{*}_{1}),206(N_{1})\end{matrix}$
$C_{4h}$ $A_{u}$ $z$ $84(A_{2},Z_{2}),85(A_{2},M_{2}),86(M_{2},Z_{2}),88(M_{2})$
$D_{2h}$ $\begin{matrix}B_{1u}\\ B_{2u}\\ B_{3u}\end{matrix}$ $z,y,x$ $\begin{matrix}51(R_{1,2},S_{1,2},U_{1,2},X_{1,2}),52(R_{1,2},Y_{1,2}),53(T_{1,2},Z_{1,2}),54(S_{1,2},X_{1,2}),\\ 55(T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),56(S_{1,2},X_{1,2},Y_{1,2}),57(S_{1,2},U_{1,2},Y_{1,2},Z_{1,2}),58(R_{1,2},X_{1,2},Y_{1,2}),\\ 59(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),60(X_{1,2},Z_{1,2}),61(X_{1,2},Y_{1,2},Z_{1,2}),\\ 62(T_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),63(T_{1,2},Z_{1,2}),64(T_{1,2},Z_{1,2}),127(R_{1,2},X_{1,2}),128(X_{1,2}),\\ 129(R_{1,2},X_{1,2}),130(X_{1,2}),135(R_{1,2},X_{1,2}),136(X_{1,2}),137(R_{1,2},X_{1,2}),\\ 138(X_{1,2}),193(L_{1,2}),194(L_{1,2}),205(X_{1,2})\end{matrix}$
$D_{4h}$ $A_{2u}$ $z$ $\begin{matrix}129(A_{3,4},M_{3,4}),130(M_{3,4}),136(A_{1,2}),137(M_{3,4}),138(M_{3,4})\end{matrix}$
$C_{2h}$ $B_{u}$ $c_{1}x+c_{2}y$ $11(C^{*}_{1},D^{*}_{1},E^{*}_{1},Z^{*}_{1}),14(C^{*}_{1},Z^{*}_{1}),63(R_{1}),176(L_{1})$
$C_{4h}$ $B_{u}$ $\begin{matrix}c_{1}xyz+\\ c_{2}z(x^{2}-y^{2})\end{matrix}$ $84(A_{1},Z_{1}),85(A_{1},M_{1}),86(M_{1},Z_{1}),88(M_{1})$
$D_{4h}$ $B_{2u}$ $z(x^{2}-y^{2})$ $\begin{matrix}126(A_{3,4}),128(A_{1,2}),129(A_{1,2},M_{1,2}),130(M_{1,2}),131(A_{1,2},Z_{1,2}),133(Z_{1,2}),135(Z_{1,2}),\\ 136(A_{3,4}),137(A_{1,2},M_{1,2},Z_{1,2}),138(M_{1,2}),141(M_{1,2}),142(M_{1,2}),223(X^{*}_{1,2})\end{matrix}$
$D_{2h}$ $A_{u}$ $xyz$ $\begin{matrix}48(S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2},Z_{1,2}),49(R_{1,2},T_{1,2},U_{1,2},Z_{1,2}),\\ 50(R_{1,2},S_{1,2},T_{1,2},U_{1,2},X_{1,2},Y_{1,2}),52(U_{1,2},X_{1,2},Z_{1,2}),53(S_{1,2},X_{1,2}),54(T_{1,2},Z_{1,2}),\\ 56(Z_{1,2}),58(Z_{1,2}),60(Y_{1,2}),66(T_{1,2},Z_{1,2}),68(T_{1,2},Z_{1,2}),70(T_{1,2},Y_{1,2},Z_{1,2}),124(R_{1,2}),\\ 125(R_{1,2},X_{1,2}),126(R_{1,2},X_{1,2}),132(R_{1,2}),133(R_{1,2},X_{1,2}),134(R_{1,2},X_{1,2}),\\ 141(X^{*}_{1,2}),142(X^{*}_{1,2}),192(L_{1,2}),201(M_{1,2},X_{1,2}),203(X_{1,2}),230(N^{*}_{1,2})\end{matrix}$
$D_{4h}$ $B_{1u}$ $xyz$ $\begin{matrix}125(A_{3,4},M_{3,4}),126(M_{3,4}),132(A_{1,2},Z_{1,2}),133(M_{1,2}),134(M_{1,2},Z_{1,2}),\\ 136(Z_{1,2}),138(Z_{1,2}),222(M_{3,4}),224(M_{1,2},X^{*}_{1,2}),227(X^{*}_{1,2}),228(X^{*}_{1,2})\end{matrix}$
$D_{4h}$ $A_{1u}$ $xyz(x^{2}-y^{2})$ $\begin{matrix}124(A_{1,2},Z_{1,2}),125(A_{1,2},M_{1,2}),126(A_{1,2},M_{1,2},Z_{1,2}),128(Z_{1,2}),130(Z_{1,2}),\\ 131(A_{3,4},Z_{3,4}),132(A_{3,4},Z_{3,4}),133(A_{1,2},M_{3,4},Z_{3,4}),134(M_{3,4},Z_{3,4}),135(Z_{3,4}),\\ 136(Z_{3,4}),137(Z_{3,4}),138(Z_{3,4}),141(M_{3,4}),142(M_{3,4}),222(M_{1,2},X^{*}_{1,2}),\\ 223(X^{*}_{3,4}),224(M_{3,4},X^{*}_{3,4}),227(X^{*}_{3,4}),228(X^{*}_{3,4})\end{matrix}$
$C_{6h}$ $B_{u}$ $\begin{matrix}c_{1}x(x^{2}-3y^{2})\\ +c_{2}y(y^{2}-3x^{2})\end{matrix}$ $176(A_{1})$
$D_{3d}$ $A_{1u}$ $x(x^{2}-3y^{2})$ $163(A_{3}),165(A_{3}),167(T_{3}),226(L^{*}_{3}),228(L^{*}_{3})$
$D_{6h}$ $A_{1u}$ $\begin{matrix}xyz(x^{2}-3y^{2})\\ \times(y^{2}-3x^{2})\end{matrix}$ $192(A_{5,6})$
$D_{6h}$ $B_{1u}$ $x(x^{2}-3y^{2})$ $194(A_{1,2})$
$D_{6h}$ $B_{2u}$ $y(y^{2}-3x^{2})$ $193(A_{1,2})$
$O_{h}$ $A_{1u}$ $\begin{matrix}xyz(x^{2}-y^{2})\times\\ (y^{2}-z^{2})(z^{2}-x^{2})\end{matrix}$ $222(R_{1}),223(R_{1}),230(H_{1})$
Table 4: Tight-binding coefficients, and odd-parity spin splittings for space groups with two atoms per unit cell related by inversion. Abbreviation $c_{i}\equiv\cos k_{i}$, $s_{i}\equiv\sin k_{i}$, $c_{i/2}\equiv\cos\frac{k_{i}}{2}$, $s_{i/2}\equiv\sin\frac{k_{i}}{2}$, $f_{x}\equiv\sin k_{x}+\sin\frac{k_{x}}{2}\cos\frac{\sqrt{3}k_{y}}{2}$, $f_{y}\equiv\sqrt{3}\cos\frac{k_{x}}{2}\sin\frac{\sqrt{3}k_{y}}{2}$, $f_{x/2}\equiv\sqrt{3}\cos\frac{k_{y}}{2\sqrt{3}}\sin\frac{k_{x}}{2}$, $f_{y/2}\equiv\sin\frac{k_{y}}{\sqrt{3}}+\sin\frac{k_{y}}{2\sqrt{3}}\cos\frac{k_{x}}{2}$, and $c_{1/2}\equiv\cos(\frac{k_{y}}{\sqrt{3}})+2\cos(\frac{k_{x}}{2})\cos(\frac{k_{y}}{2\sqrt{3}})$ applies. Wave-vectors $\bf Q$ that do not host 2D irreducible representations are not included. $(s_{x},s_{z})$ denotes linear combination of $s_{x}$ and $s_{z}$. Only leading hopping coefficients are included, while the spin splitting can come from higher-order hopping coefficients. For example, $s_{x}s_{y}s_{z}$ in 11(2e) requires $t_{x}=c_{1}c_{y/2}+c_{2}s_{x}c_{y/2}s_{z}$. Trivial factors like $c_{x}+c_{y}$ have been omitted.
SG $t_{x}$ $t_{y}$ $\begin{pmatrix}\pi\\ 0\\ 0\end{pmatrix}$ $\begin{pmatrix}0\\ \pi\\ 0\end{pmatrix}$ $\begin{pmatrix}0\\ 0\\ \pi\end{pmatrix}$ $\begin{pmatrix}0\\ \pi\\ \pi\end{pmatrix}$ $\begin{pmatrix}\pi\\ 0\\ \pi\end{pmatrix}$ $\begin{pmatrix}\pi\\ \pi\\ 0\end{pmatrix}$ $\begin{pmatrix}\pi\\ \pi\\ \pi\end{pmatrix}$
11(2e) $c_{y/2}$ $c_{y/2}(s_{x},s_{z})$ $\begin{matrix}s_{y},\\ s_{x}s_{y}s_{z}\end{matrix}$ $\begin{matrix}s_{y},\\ s_{x}s_{y}s_{z}\end{matrix}$ $\begin{matrix}s_{y},\\ s_{x}s_{y}s_{z}\end{matrix}$ $\begin{matrix}s_{y},\\ s_{x}s_{y}s_{z}\end{matrix}$
13(2e,2f) $c_{z/2}$ $s_{y}(c_{z/2},s_{x}s_{z/2})$ $s_{x},s_{z}$ $s_{x},s_{z}$ $s_{x},s_{z}$ $s_{x},s_{z}$
48(2a-2d) $c_{x/2}c_{y/2}c_{z/2}$ $s_{x/2}s_{y/2}s_{z/2}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}$ $s_{y}$ $s_{z}$
49(2e-2h) $c_{z/2}$ $s_{x}s_{y}s_{z/2}$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
50(2a-2d) $c_{x/2}c_{y/2}$ $s_{x/2}s_{y/2}s_{z}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$ $s_{z}$ $s_{z}$
51(2e,2f) $c_{x/2}$ $c_{x/2}s_{z}$ $s_{x}$ $s_{x}$ $s_{x}$ $s_{x}$
59(2a,2b) $c_{x/2}c_{y/2}$ $c_{x/2}c_{y/2}s_{z}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$ $s_{x}s_{y}s_{z}$ $s_{x}s_{y}s_{z}$
84(2e,2f) $c_{z/2}$ $\begin{matrix}s_{x}s_{y}s_{z/2},\\ (c_{x}-c_{y})s_{z/2}\end{matrix}$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
85(2a,2b) $c_{x/2}c_{y/2}$ $\begin{matrix}s_{x/2}s_{y/2}s_{z},\\ c_{x/2}c_{y/2}(c_{x}-c_{y})s_{z}\end{matrix}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$ $\begin{matrix}s_{z},\\ (c_{x}-c_{y})s_{x}s_{y}s_{z}\end{matrix}$ $\begin{matrix}s_{z},\\ (c_{x}-c_{y})s_{x}s_{y}s_{z}\end{matrix}$
85(2c) $c_{x/2}c_{y/2}$ $\begin{matrix}c_{x/2}c_{y/2}s_{z},\\ s_{x/2}s_{y/2}(c_{x}-c_{y})s_{z}\end{matrix}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$ $\begin{matrix}s_{x}s_{y}s_{z},\\ (c_{x}-c_{y})s_{z}\end{matrix}$ $\begin{matrix}s_{x}s_{y}s_{z},\\ (c_{x}-c_{y})s_{z}\end{matrix}$
86(2a,2b) $c_{x/2}c_{y/2}c_{z/2}$ $\begin{matrix}s_{x/2}s_{y/2}s_{z/2},\\ c_{x/2}c_{y/2}(c_{x}-c_{y})s_{z/2}\end{matrix}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x},s_{y}$ $s_{x},s_{y}$ $\begin{matrix}s_{z},\\ (c_{x}-c_{y})s_{x}s_{y}s_{z}\end{matrix}$
124(2a,2c) $c_{z/2}$ $s_{x}s_{y}(c_{x}-c_{y})s_{z/2}$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
125(2a,2b) $c_{x/2}c_{y/2}$ $s_{x/2}s_{y/2}(c_{x}-c_{y})s_{z}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$ $(c_{x}-c_{y})s_{z}$ $(c_{x}-c_{y})s_{z}$
125(2c,2d) $c_{x/2}c_{y/2}$ $s_{x/2}s_{y/2}s_{z}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$ $s_{z}$ $s_{z}$
126(2a,2b) $c_{x/2}c_{y/2}c_{z/2}$ $s_{x/2}s_{y/2}(c_{x}-c_{y})s_{z/2}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}$ $s_{y}$ $(c_{x}-c_{y})s_{z}$ $s_{x}s_{y}s_{z}$
129(2a,2b) $c_{x/2}c_{y/2}$ $c_{x/2}c_{y/2}(c_{x}-c_{y})s_{z}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$ $s_{x}s_{y}(c_{x}-c_{y})s_{z}$ $s_{x}s_{y}(c_{x}-c_{y})s_{z}$
129(2c) $c_{x/2}c_{y/2}$ $c_{x/2}c_{y/2}s_{z}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$ $s_{x}s_{y}s_{z}$ $s_{x}s_{y}s_{z}$
131(2e,2f) $c_{z/2}$ $(c_{x}-c_{y})s_{z/2}$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
132(2b,2d) $c_{z/2}$ $s_{x}s_{y}s_{z/2}$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
134(2a,2b) $c_{x/2}c_{y/2}c_{z/2}$ $s_{x/2}s_{y/2}s_{z/2}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}$ $s_{y}$ $s_{z}$
137(2a,2b) $c_{x/2}c_{y/2}c_{z/2}$ $c_{x/2}c_{y/2}(c_{x}-c_{y})s_{z/2}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{y}$ $s_{x}$ $s_{x}s_{y}(c_{x}-c_{y})s_{z}$ $s_{x}s_{y}s_{z}$
201(2a) $c_{x/2}c_{y/2}c_{z/2}$ $s_{x/2}s_{y/2}s_{z/2}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}$ $s_{y}$ $s_{z}$
222(2a) $c_{x/2}c_{y/2}c_{z/2}$ $\begin{matrix}s_{x/2}s_{y/2}s_{z/2}(c_{y}-c_{z})\\ (c_{z}-c_{x})(c_{x}-c_{y})\end{matrix}$ $s_{x}$ $s_{y}$ $s_{z}$ $\begin{matrix}s_{x}\times\\ (c_{y}-c_{z})\end{matrix}$ $\begin{matrix}s_{y}\times\\ (c_{z}-c_{x})\end{matrix}$ $\begin{matrix}s_{z}\times\\ (c_{x}-c_{y})\end{matrix}$ $s_{x}s_{y}s_{z}$
224(2a) $c_{x/2}c_{y/2}c_{z/2}$ $s_{x/2}s_{y/2}s_{z/2}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}$ $s_{y}$ $s_{z}$
SG $t_{x}$ $t_{y}$ $\begin{pmatrix}0\\ 0\\ \pi\end{pmatrix}$ $\begin{pmatrix}\pi\\ 0\\ 0\end{pmatrix}$ $\begin{pmatrix}\pi\\ 0\\ \pi\end{pmatrix}$
163(2a) $c_{z/2}$ $f_{y}(f_{y}^{2}-3f_{x}^{2})c_{z/2}$ $s_{z}$ $s_{z}$
163(2c,2d) $c_{1/2}c_{z/2}$ $f_{y/2}(f_{y/2}^{2}-3f_{x/2}^{2})c_{z/2}$ $s_{z}$ $s_{z}$
165(2a) $c_{z/2}$ $f_{x}(f_{x}^{2}-3f_{y}^{2})c_{z/2}$ $s_{z}$ $s_{z}$
176(2a) $c_{z/2}$ $\begin{matrix}f_{x}(f_{x}^{2}-3f_{y}^{2})c_{z/2},\\ f_{y}(f_{y}^{2}-3f_{x}^{2})c_{z/2}\end{matrix}$ $s_{z}$ $s_{z}$
176(2c,2d) $c_{1/2}c_{z/2}$ $\begin{matrix}f_{y/2}(f_{y/2}^{2}-3f_{x/2}^{2})c_{z/2},\\ f_{x/2}(f_{x/2}^{2}-3f_{y/2}^{2})c_{z/2}\end{matrix}$ $s_{z}$ $s_{z}$
192(2a) $c_{z/2}$ $\begin{matrix}f_{x}(f_{x}^{2}-3f_{y}^{2})\times\\ f_{y}(f_{y}^{2}-3f_{x}^{2})s_{z/2}\end{matrix}$ $s_{z}$ $s_{z}$
193(2a) $c_{z/2}$ $f_{x}(f_{x}^{2}-3f_{y}^{2})c_{z/2}$ $s_{z}$ $s_{z}$
194(2b) $c_{z/2}$ $f_{y}(f_{y}^{2}-3f_{x}^{2})c_{z/2}$ $s_{z}$ $s_{z}$
194(2c,2d) $c_{1/2}c_{z/2}$ $f_{y/2}(f_{y/2}^{2}-3f_{x/2}^{2})c_{z/2}$ $s_{z}$ $s_{z}$
Table 5: Tight-binding coefficients, and odd-parity spin splittings for space groups with two atoms per unit cell at inversion center. The same abbreviation applies.
SG $t_{x}$ $t_{z}$ $\begin{pmatrix}\pi\\ 0\\ 0\end{pmatrix}$ $\begin{pmatrix}0\\ \pi\\ 0\end{pmatrix}$ $\begin{pmatrix}0\\ 0\\ \pi\end{pmatrix}$ $\begin{pmatrix}0\\ \pi\\ \pi\end{pmatrix}$ $\begin{pmatrix}\pi\\ 0\\ \pi\end{pmatrix}$ $\begin{pmatrix}\pi\\ \pi\\ 0\end{pmatrix}$ $\begin{pmatrix}\pi\\ \pi\\ \pi\end{pmatrix}$
11(2a-2d) $c_{y/2}$ $s_{y}(s_{x},s_{z})$ $s_{y}$ $s_{y}$ $s_{y}$ $s_{y}$
13(2a-2d) $c_{z/2},s_{x}s_{z/2}$ $s_{y}(s_{x},s_{z})$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
14(2a-2d) $c_{y/2}(s_{x}s_{z/2},c_{z/2})$ $s_{y}(s_{x},s_{z})$ $s_{y}$ $s_{z}$ $s_{z}$ $s_{y}$
49(2a-2d) $c_{z/2}$ $s_{x}s_{y}$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
51(2a-2d) $c_{x/2}$ $s_{x}s_{z}$ $s_{x}$ $s_{x}$ $s_{x}$ $s_{x}$
53(2a-2d) $c_{x/2}c_{z/2}$ $s_{y}s_{z}$ $s_{x}$ $s_{z}$ $s_{z}$ $s_{x}$
55(2a-2d) $c_{x/2}c_{y/2}$ $s_{x}s_{y}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$
58(2a-2d) $c_{x/2}c_{y/2}c_{z/2}$ $s_{x}s_{y}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}s_{y}s_{z}$
84(2a,2b) $c_{z/2}$ $(c_{x}-c_{y}),s_{x}s_{y}$ $s_{z}$ $s_{z}$
84(2c,2d) $\begin{matrix}c_{x/2}c_{y/2}c_{z/2},\\ s_{x/2}s_{y/2}c_{z/2}\end{matrix}$ $(c_{x}-c_{y}),s_{x}s_{y}$ $s_{z}$ $s_{x}s_{y}s_{z}$
124(2b,2d) $c_{z/2}$ $s_{x}s_{y}(c_{x}-c_{y})$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
127(2a,2b) $c_{x/2}c_{y/2}$ $s_{x}s_{y}(c_{x}-c_{y})$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$
127(2c,2d) $c_{x/2}c_{y/2}$ $s_{x}s_{y}$ $s_{x}$ $s_{y}$ $s_{y}$ $s_{x}$
128(2a,2b) $c_{x/2}c_{y/2}c_{z/2}$ $s_{x}s_{y}(c_{x}-c_{y})$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}s_{y}s_{z}$
131(2a,2b) $c_{z/2}$ $(c_{x}-c_{y})$ $s_{z}$ $s_{z}$
131(2c,2d) $c_{x/2}c_{y/2}c_{z/2}$ $(c_{x}-c_{y})$ $s_{z}$ $s_{x}s_{y}s_{z}$
132(2a,2c) $c_{z/2}$ $s_{x}s_{y}$ $s_{z}$ $s_{z}$ $s_{z}$ $s_{z}$
136(2a,2b) $c_{x/2}c_{y/2}c_{z/2}$ $s_{x}s_{y}$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}s_{y}s_{z}$
223(2a) $c_{x/2}c_{y/2}c_{z/2}$ $(c_{x}-c_{y})(c_{y}-c_{z})(c_{z}-c_{x})$ $s_{x}$ $s_{y}$ $s_{z}$ $s_{x}s_{y}s_{z}$
SG $t_{x}$ $t_{z}$ $\begin{pmatrix}0\\ 0\\ \pi\end{pmatrix}$ $\begin{pmatrix}\pi\\ 0\\ 0\end{pmatrix}$ $\begin{pmatrix}\pi\\ 0\\ \pi\end{pmatrix}$
163(2b) $\begin{matrix}c_{z/2},\\ f_{x}(3f_{y}^{2}-f_{x}^{2})s_{z/2}\end{matrix}$ $\begin{matrix}f_{y}(f_{y}^{2}-3f_{x}^{2})s_{z},\\ f_{x}f_{y}(f_{x}^{2}-3f_{y}^{2})(3f_{x}^{2}-f_{y}^{2})\end{matrix}$ $s_{z}$ $s_{z}$
165(2b) $\begin{matrix}c_{z/2},\\ f_{y}(3f_{x}^{2}-f_{y}^{2})s_{z/2}\end{matrix}$ $\begin{matrix}f_{x}(f_{x}^{2}-3f_{y}^{2})s_{z},\\ f_{x}f_{y}(f_{x}^{2}-3f_{y}^{2})(3f_{x}^{2}-f_{y}^{2})\end{matrix}$ $s_{z}$ $s_{z}$
176(2b) $c_{z/2}$ $\begin{matrix}s_{z}f_{x}(f_{x}^{2}-3f_{y}^{2}),\\ s_{z}f_{y}(f_{y}^{2}-3f_{x}^{2})\end{matrix}$ $s_{z}$ $s_{z}$
192(2b) $c_{z/2}$ $f_{x}f_{y}(f_{x}^{2}-3f_{y}^{2})(3f_{x}^{2}-f_{y}^{2})$ $s_{z}$ $s_{z}$
193(2b) $c_{z/2}$ $s_{z}f_{x}(f_{x}^{2}-3f_{y}^{2})$ $s_{z}$ $s_{z}$
194(2a) $c_{z/2}$ $s_{z}f_{y}(f_{y}^{2}-3f_{x}^{2})$ $s_{z}$ $s_{z}$
Appendix F Electronic structure of FeSe
Density-functional theory calculations for bulk FeSe with the $(\pi,\pi,0)$ coplanar magnetic order are carried out using the Full-potential Linearized Augmented Plane Wave (FLAPW) method Weinert et al. (2009), with the scalar-relativistic approximation Koelling and Harmon (1977); MacDonald et al. (1980). The Perdew-Burke-Ernzerhof form of exchange correlation functional Perdew et al. (1996) is used, along with wave function and potential energy cutoffs of 16 and 200 Ry, respectively. A muffin-tin sphere radius of 1.1 Å is applied to both Fe and Se atoms. The superlattice setup is $\sqrt{2}\times\sqrt{2}\times 1$, as indicated by the dotted line in the left panel of Figure 2 of the main text, with experimental lattice parameters ($a=3.765$; $c=5.518$; $z_{\textrm{Se}}=0.26$) Villars (1997). The Brillouin zone is sampled using (15,15,10) and (100,100,50) k-point meshes for self-consistent field calculation and Fermi surface calculation, respectively. The resulting band structure agrees with a previous calculation employing the generalized Bloch theorem for spin spirals Shishidou et al. (2018).
Appendix G Electronic structure of CeNiAsO
For the construction of the minimal low-energy models relevant for CeNiAsO, we perform ab initio calculations within the full-potential local-orbital (FPLO) code Koepernik and Eschrig (1999) (version 22.00-62) in the non-relativistic setting and the LDSA approximation. We use $a=b=4.014537\mathrm{\AA}$, $c=7.90723\mathrm{\AA}$ in space group 129, Ce on 2c with internal coordinates $(1/4,1/4,0.154739)$, Ni on 2b with $(1/4,1/4,1/2)$, As on 2c with $(1/4,1/4,-0.351368)$ and O on 2a with $(1/4,3/4,0)$. The resulting band structure is shown in Fig. 5 which exhibits two pairs of not hybridized bands as checked by calculating the irreps of the corresponding bands.
Figure 5: Electronic structure of CeNiAsO including orbital projections to Ce and Ni states. Two pairs of bands cross the Fermi level $E_{F}$ such that two single band models are relevant to describe the low-energy electronic structure.